# Joint Score-Threshold Optimization for Interpretable Risk Assessment Under Partial Supervision

### v1

@@ 5 Discussion
5.1 Key Findings
The proposed CSO and mixed-integer programming pipeline achives significant performance gains compared to the current JHFRAT model while offering priority-based tuning options in the hyperparameters. In general, the score distributions and sample categorizations respond as expected to the provided hyperparameters. Utilizing a symmetric distance function for mis-classification loss results in nearly symmetrical mis-classification rates for low-risk and high-risk labeled samples, while the use of an asymmetric 3-to-1 high-risk to low-risk distance function results in higher skew of both classes towards high-risk categorization. The notable exception to expectations is the increase in low-risk false negatives among the two performance-constrained models. The false-negative constraint for the FN&FP-cap MIP should mitigate an increase in such classifications, but this cannot be guaranteed in the testing setting.
While the CSO and MIP models themselves do not inherently discourage moderate-risk classification, when applied in a setting with no such moderate risk labels, the resulting score distributions disaggregate from the mean. This movement away from moderate risk classification can come at the expense of increasing further-class mis-classification, as is the case for the FP-cap and FN&FP-cap models. Nevertheless, these models remain promising options for improving true high-risk classification without drastic increases in the total number of patients classified as high-risk.
5.2 Limitations and Future Work
Several model extensions warrant investigation:
  • Non-linear scoring functions: Extending to tree-based or shallow neural architectures while maintaining interpretability
  • Time-varying risk: Incorporating temporal dynamics in sequential assessment settings
  • Multi-site learning: Federated optimization across institutions with heterogeneous populations
There are a few key limitations with the JHFRAT model experiments. Use of average JHFRAT scores over the encounter for model training results in some non-binary risk score values. These binary risk-score values represent averages over the encounter, but may contribute to extreme values in the optimized coefficients (i.e. coefficients larger than the highest class lower bound) in the absence of enforced upper bounds. The exclusion of all intervention cohort encounters also potentially limits the generalizability of the optimized models. Future work could explore including some patients from the intervention cohort as Safe Moderate-Risk, or apply multiple labels, depending on the number and type of fall prevention interventions applied.
5.3 Clinical Deployment Considerations
It is of particular clinical interest to improve identification of patients at high risk for falls while avoiding an overall increase in risk scores. From a resource perspective, it is impractical and wasteful to provide fall prevention interventions to patients who do not need them. Furthermore, from a clinical perspective, restricting mobility is an effective fall prevention method, but often leads to detrimental functional decline. Therefore, placing such restrictions on low-risk patients is not only unnecessary, but is actively harmful. These types of concerns over false positives extend beyond fall risk assessment. Alert fatigue is a persistent problem that is exacerbated by many false positives in assessment of risks of adverse events such as sepsis. The inclusion of performance constraints and the user-tunable distance function for mis-classification loss serve as tools to control the balance between under-classification and over-classification. Specific constraints on the number of false positives and/or false negatives for the most extreme classes offer assuredness.
In general, the specific features of the proposed MIP formulation that are applied to any given setting can be based on the context of the risk factors, adverse event in question, available data, and current clinical practice. For example, enforcing optimized coefficients to be integers may be unnecessary if assessments are completed automatically and clinicians only observe the final score. Allowing negativity in model coefficients allows for the inclusion of risk-preventative factors in the model.
@@ 6 Conclusion
In this work we develop a multi-class mixed integer programming ordinal classification model and a smooth, convex constrained score optimization. These models can apply in a variety of risk assessment contexts, and are particularly well suited for settings with uncertain ground truth or predictive performance requirements compared to an existing model. We utilize the proposed MIP and CSO formulations to conduct data-driven score optimization of the Johns Hopkins Fall Risk Assessment Tool. Across different model variations, we are able to increase identification of patients who fell as high-risk, and control for false positives via model parameters and constraints. Future work can explore refining or adapting the proposed models to other risk assessment settings.

### v2

@@ 6 Discussion
The proposed framework makes several methodological contributions to the optimization of clinical risk scoring tools. First, the formulation addresses a common challenge in clinical settings where reliable labels exist only for extreme risk categories. When optimizing with only these extreme labels, the model may collapse intermediate thresholds to eliminate unlabeled middle categories, effectively producing a binary classifier rather than the desired ordinal tool. The framework prevents this degeneracy through minimum threshold gap constraints, maintaining clinically meaningful risk stratification without requiring labeled examples for every category. Second, the asymmetric distance-aware loss function directly encodes clinical priorities into the optimization objective. By allowing differential penalties for under-triage versus over-triage that scale with ordinal distance, the framework aligns the training signal with deployment consequences, which is preferable to learning symmetric costs and adjusting thresholds post-hoc. Third, the governance and performance constraints provide mechanisms for encoding domain knowledge and ensuring operational safety, from sign restrictions that maintain clinical face validity to false-positive and false-negative caps that guarantee minimum performance levels. Finally, the two-phase optimization pipeline combines the computational efficiency of convex optimization with the exactness of mixed-integer programming, enabling tractable solutions for moderate-sized clinical datasets.
The JHFRAT case study validates these methodological contributions empirically. Training only on the 8,348 patients with confident extreme labels while excluding the 38,643 with intervention-censored outcomes yields models that generalize well to held-out test data, with tight accuracy improving from 69% to 82% and high-risk recall nearly doubling from 29% to 48%. The asymmetric loss function (3:1 under-triage to over-triage penalty) successfully biases the optimization toward high-risk identification without sacrificing overall discrimination, as evidenced by the AUPRC improvement from 0.51 to 0.67. The two-phase pipeline demonstrates practical value: the CSO warm-start achieves 79% tight accuracy on its own and provides feasible integer solutions that accelerate MIP convergence, while the MIP refinement yields additional gains in threshold-based classification. The similar AUC scores but improved accuracy metrics between CSO and MIP suggest that while both methods learn comparable score rankings, the MIP’s discrete optimization produces coefficients better calibrated to the fixed threshold boundaries.
Beyond performance improvements, the optimized coefficients reveal clinically meaningful patterns that illustrate the value of data-driven calibration. Cognition-related factors, particularly altered awareness and impulsivity, receive substantially higher weights in the optimized model, increasing from 1–2 points to 12–13 points. Elimination-related factors similarly receive increased weights, aligning with research on bathroom-related ambulation as a high-risk period. Conversely, patient care equipment coefficients are reduced to zero, possibly reflecting confounding in the original design where equipment presence correlates with closer monitoring or serves as a proxy for illness severity captured elsewhere. These discoveries would not emerge from simply re-thresholding the original scores; they require the coefficient optimization that the framework enables.
Several limitations warrant acknowledgment. The use of encounter-averaged JHFRAT scores introduces non-integer feature values, which may contribute to the large coefficient magnitudes observed; future work could optimize directly on assessment-level data or impose tighter coefficient bounds. The complete exclusion of the intervention cohort, while methodologically principled, may limit generalizability to patients who would typically receive interventions. Incorporating partial labels based on the specific interventions applied or otherwise implementing multi-label formulations could address this. The fixed-threshold constraint, while practically motivated for deployment compatibility, prevents full exploitation of the joint optimization framework. Additionally, the widened score range (maximum scores increasing from 33 to 56) may affect clinician interpretation of within-category gradation, requiring communication and training for deployment.
The framework applies broadly to clinical risk stratification settings sharing the key structural features: ordinal categories from point-based scores, labels available only at extreme categories, and asymmetric misclassification costs. Examples include pressure injury risk (Braden Scale), venous thromboembolism risk (Caprini Score), and early warning scores for deterioration. Future extensions include non-linear scoring functions maintaining interpretability, temporal dynamics for time-varying risk, and federated optimization across institutions with heterogeneous populations.
@@ 7 Conclusion
This work presents a mixed-integer programming framework for optimizing clinical risk scoring tools in settings where reliable labels exist only for extreme risk categories. The framework addresses three challenges that impede standard supervised learning in clinical deployment: missing labels for intermediate categories that can cause threshold collapse, asymmetric misclassification costs that vary with ordinal distance, and interpretability requirements that mandate integer coefficients amenable to manual calculation. Minimum threshold gap constraints maintain clinically meaningful risk stratification without requiring labeled examples for every category, while the constrained score optimization relaxation provides an efficient warm-start strategy that accelerates MIP convergence.
Applied to the Johns Hopkins Fall Risk Assessment Tool, the optimized model improves high-risk recall from 29% to 48% and area under the precision-recall curve from 0.51 to 0.67, while maintaining the existing three-tier threshold structure for seamless integration with current clinical workflows. The optimized coefficients reveal that cognition and elimination factors are substantially more predictive than their original weights suggest, while patient care equipment contributes minimal independent information. These insights may inform future tool revisions beyond the current optimization.
The framework’s flexibility in accommodating governance constraints, performance guarantees, and minimal-edit requirements positions it for broader application across clinical risk assessment domains where data-driven calibration of legacy tools is needed. Future extensions may incorporate time-varying risk dynamics, multi-site federated learning, and richer label structures that leverage intervention cohort information more directly.